# Student Perspectives on the Benefits and Risks of AI in Education

### v1

@@ 3.2.2 Accuracy of Information
Information reliability constituted the second most significant concern. 91 statements were coded in this category, 26.38% of the total. Participants expressed wariness about misinformation, hallucinated data, and the absence of source citations in AI responses. One respondent stated that their main concern was “the chatbot giving incorrect and fabricated information. If a student were to use one, and were given false information, they would not know it was false unless they put in the effort to cross reference it.” This reflects anxiety not only about AI-generated inaccuracies but also about students’ potential failure to verify such information, even if they are aware of the systems not being fully reliable. One respondent noted that AI chatbots “do not always give accurate information as they are not ”aware” of what they are discussing. They are simply giving the best response to a series of words.” Despite being aware of this, many respondents were concerned that students might be misled by inaccurate responses.
@@ 3.2.3 Loss of Critical Thinking
Another major concern for respondents was students’ loss of critical thinking skills when using AI chatbots. This theme was found in 67 statements, 19.42% of the total. Many respondents noted concerns of students’ no longer putting effort into properly learning information or completing assignments, simply having AI generate a response for them. Respondents specifically noted that students might use conversation AI to “avoid learning material” by asking chatbots for answers “just . . . to complete an assignment, not to learn,” and “it is very easy for students to easily find the answers without needing to actually understand the content.” One respondent even noted that “students can pass hard courses [by] putting minimal effort and … using various AI tools. The student will not gain knowledge but will only cheat themselves.”
@@ 4 Discussion
The largest concern students identified with using AI chatbots in education was academic integrity. The emphasis on academic integrity aligns with prior studies [17, 29, 22, 20, 30] that identified plagiarism concerns. However, students worry not only about intentional cheating but also about their legitimate work being mistakenly flagged as AI-generated, presenting a more complex challenge for instructors. The second most cited concern was accuracy of information. Unlike academic integrity, where the focus was on misuse, students emphasized concerns over the reliability of AI-generated information and assistance, aligning with concerns found in previous research [24, 27, 31, 32, 7]. Students worried about “the chatbot giving incorrect and fabricated information,” and that without proper verification, students “would not know it was false unless they put in the effort to cross reference it.” The third most cited concern was loss of critical thinking skills. Students worried about peers “avoiding learning material” by relying too heavily on AI, with which students might “pass hard courses [by] putting minimal effort” while not gaining actual knowledge. The concerns about skill degradation and overreliance align with and extend prior research [17, 24, 22, 30], as students specifically responded with concerns to their diminishing ability to write and think independently.
@@ 4.1 Limitations and Future Work
While several insights are provided into student perspectives on AI chatbots in education, the limitations of this study suggest directions for future research. This study focused on undergraduate students at a limited number of institutions. Future studies could assess perspectives across different types of institutions, academic levels, and geographical regions to develop a more comprehensive understanding of student perspectives. Second, while this study found student concerns relating to potential loss of critical thinking skills and overreliance, understanding the long-term impacts of AI chatbots on learning outcomes requires additional research. Future studies could examine how different patterns of AI chatbot usage affect student performance, skill development, and learning retention. Third, attitudes toward AI chatbots captured in this study represent a snapshop during a period of rapid technological advancement and shifing social acceptance of AI. Future longitudinal studies can examine how student perceptions change alongside technological advancements, changing institutional policies, and broader societal discourse around AI in education. Such research could help identify whether concerns about academic integrity and information accuracy persist as technologies mature and institutional policies adapt.
@@ 5 Conclusion
The findings demonstrate alignment between student perspectives and prior research demonstrating potential the benefits of AI in education. Students valued the feedback and study support, instruction capabilities, increased access to information, increased productivity, increased creativity offered by AI chatbots.They specifically appreciated immediate assistance when instructors were unavailable, help with organizing tasks and brainstorming ideas, and clear explanations of complex topics. However, students also expressed concerns that both paralleled and expanded upon previous research. Academic integrity concerns extended beyond plagiarism to include worries about legitimate work being misidentified as AI-generated. Concerns additionally related to the accuracy and reliability of information, the loss of critical thinking and writing skills, a the potential development of overreliance. Students additionally raised ethical considerations regarding data privacy, system bias, environmental impact, and the preservation of human elements in education.

### v2

@@ 3.2.2 Accuracy of Information
Information reliability constituted the second most significant concern. 91 statements were coded in this category, 26.38% of the total. Participants expressed wariness about misinformation, hallucinated data, and the absence of source citations in AI responses. One respondent stated that their main concern was “the chatbot giving incorrect and fabricated information. If a student were to use one, and were given false information, they would not know it was false unless they put in the effort to cross reference it.” This reflects anxiety not only about AI-generated inaccuracies but also about students potential failure to verify such information, even if they are aware of the systems not being fully reliable. One respondent noted that AI chatbots “do not always give accurate information as they are not ”aware” of what they are discussing. They are simply giving the best response to a series of words.” Despite being aware of this, many respondents were concerned that students might be misled by inaccurate responses.
@@ 3.2.3 Loss of Critical Thinking
Another major concern for respondents was students loss of critical thinking skills when using AI chatbots. This theme was found in 67 statements, 19.42% of the total. Many respondents noted concerns of students no longer putting effort into properly learning information or completing assignments, simply having AI generate a response for them. Respondents specifically noted that students might use conversation AI to “avoid learning material” by asking chatbots for answers “just . . . to complete an assignment, not to learn,” and “it is very easy for students to easily find the answers without needing to actually understand the content.” One respondent even noted that “students can pass hard courses [by] putting minimal effort and … using various AI tools. The student will not gain knowledge but will only cheat themselves.”
@@ 4 Discussion
The largest concern students identified with using AI chatbots in education was academic integrity. The emphasis on academic integrity aligns with prior studies [17, 29, 22, 20, 30] that identified plagiarism concerns. However, students worry not only about intentional cheating but also about their legitimate work being mistakenly flagged as AI-generated, presenting a more complex challenge for instructors. The second most cited concern was accuracy of information. Unlike academic integrity, where the focus was on misuse, students emphasized concerns over the reliability of AI-generated information and assistance, aligning with concerns found in previous research [24, 27, 31, 32, 7]. Students worried about “the chatbot giving incorrect and fabricated information,” and that without proper verification, students “would not know it was false unless they put in the effort to cross reference it.” The third most cited concern was loss of critical thinking skills. Students noted peers “avoiding learning material” by relying too heavily on AI, with which students might “pass hard courses [by] putting minimal effort” while not gaining actual knowledge. The concerns about skill degradation and overreliance align with and extend prior research [17, 24, 22, 30], as students specifically responded with concerns to their diminishing ability to write and think independently.
@@ 4.1 Limitations and Future Work
While several insights are provided into student perspectives on AI chatbots in education, the limitations of this study suggest directions for future research. This study focused on undergraduate students at a limited number of institutions. Future studies could assess perspectives across different types of institutions, academic levels, and geographical regions to develop a more comprehensive understanding of student perspectives. Second, while this study found student concerns relating to potential loss of critical thinking skills and overreliance, understanding the long-term impacts of AI chatbots on learning outcomes requires additional research. Future studies could examine how different patterns of AI usage affect student performance, skill development, and learning retention. Third, attitudes toward AI chatbots captured in this study represent a snapshot during a period of rapid technological advancement and shifting social acceptance of AI. Future longitudinal studies can examine how student perceptions change alongside technological advancements, changing institutional policies, and broader societal discourse around AI in education. Such research could help identify whether concerns about academic integrity and information accuracy persist as technologies mature and institutional policies adapt.
@@ 5 Conclusion
The findings demonstrate alignment between student perspectives and prior research demonstrating potential the benefits of AI in education. Students valued the feedback and study support, instruction capabilities, increased access to information, increased productivity, increased creativity offered by AI chatbots. They specifically appreciated immediate assistance when instructors were unavailable, help with organizing tasks and brainstorming ideas, and clear explanations of complex topics. Students’ concerns paralleled and expanded upon previous research. Academic integrity concerns extended beyond plagiarism to include worries about legitimate work being misidentified as AI-generated. Additional concerns related to the accuracy and reliability of information, the loss of critical thinking and writing skills, the potential development of overreliance, and ethical considerations regarding data privacy, system bias, environmental impact, and the preservation of human elements in education.